### v1

# thanks: Authors contributed equallythanks: Authors contributed equallythanks: Authors contributed equally

Abstract
Quantum error correction (QEC) provides a practical path to fault-tolerant quantum computing through scaling to large qubit numbers, assuming that physical errors are sufficiently uncorrelated in time and space. In superconducting qubit arrays, high-energy impact events produce correlated errors, violating this key assumption. Following such an event, phonons with energy above the superconducting gap propagate throughout the device substrate, which in turn generate a temporary surge in quasiparticle (QP) density throughout the array. When these QPs tunnel across the qubits’ Josephson junctions, they induce correlated errors. Engineering different superconducting gaps across the qubit’s Josephson junctions provides a method to resist this form of QP tunneling. By fabricating all-aluminum transmon qubits with both strong and weak gap engineering on the same substrate, we observe starkly different responses during high-energy impact events. Strongly gap engineered qubits do not show any degradation in $T_{1}$ during impact events, while weakly gap engineered qubits show events of correlated degradation in $T_{1}$. We also show that strongly gap engineered qubits are robust to QP poisoning from increasing optical illumination intensity, whereas weakly gap engineered qubits display rapid degradation in coherence. Based on these results, gap engineering removes the threat of high-energy impacts to QEC in superconducting qubit arrays.
Scalable and useful quantum computing is achievable through the use of quantum error correction (QEC), provided the hardware satisfies key underlying assumptions. One such assumption is the absence of any substantial source of correlated error. In superconducting qubit arrays, this assumption is heavily violated by error bursts due to high-energy impacts [1, 2, 3, 4, 5, 6, 7]. During an impact event, high-energy radiation deposits up to 1 MeV in the device substrate, which radiates throughout the device as high-energy phonons [1, 8]. As illustrated in Fig. 1(a), these phonons then break Cooper pairs, producing significantly elevated quasiparticle (QP) densities. These QPs tunnel through the Josephson junction, absorbing the qubit excitation and inducing excess $T_{1}$ decay [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Impacts therefore produce a massive correlated error burst that extends through time and space, which is not correctable using standard QEC. The presence of these bursts induces an error floor on the logical performance of QEC experiments, preventing the promised exponential suppression of logical error with increasing code size [20, 21]. Without mitigation in hardware, this effect single-handedly prevents scaling QEC to algorithmically relevant logical error rates.
Figure 1: Suppressing QP tunneling using gap engineering. a) High-energy impact events deposit large energies in the device substrate, which radiates as high-energy phonons. These phonons break Cooper pairs, generating QPs which tunnel through the Josephson junction and induce $T_{1}$ decay. b) QP tunneling by absorbing the qubit energy is only possible when the qubit energy exceeds the difference in superconducting gaps between the two junction leads. c) The superconducting gap of aluminum is sensitive to film thickness, permitting access to different regimes of gap difference. d) The device consists of a checkerboard pattern of weakly (orange) and strongly (green) gap engineered transmon qubits. A device diagram to scale is included in the Supplementary Material.
While there have been several suggested [22] and implemented strategies for suppressing these effects, none have yet demonstrated total suppression of the error bursts. Moving to a low-radiation underground facility [2, 23] greatly reduces the event rate of impacts, but remaining events such as those due to radioactive materials in the device and packaging remain unmitigated [5]. Using phonon or QP traps [24, 25, 4, 26] reduces both the temporal and spatial extent of the error burst, but QP poisoning is still observed.
Meanwhile, gap engineering has been shown to modify the dynamics of QP tunneling and improve the charge-parity lifetimes in superconducting qubits [9, 10, 27, 28, 29, 30, 31, 32, 33, 34, 35]. It has also been shown to improve the speed of recovery from QP poisoning from high-energy impact events [6]. Our approach to gap engineering involves producing different superconducting gaps on each side of the junction presents an energy barrier to QPs, as illustrated in Fig. 1(b). The superconducting gap in thin films of aluminum is sensitive to thickness, with thinner films displaying higher superconducting gaps. When the gap difference between the thin and thick leads of the Josephson junction ($\delta\Delta=\Delta_{\mathrm{thin}}-\Delta_{\mathrm{thick}}$) is significantly larger than the qubit energy ($E_{q}=hf_{q}$), QPs with energy near the gap can no longer absorb the qubit energy required to climb up the barrier to the higher gap material. This suppresses the primary mechanism by which QPs induce $T_{1}$ decay.
Here, we investigate the prevalence of correlated error bursts on a single device with two types of transmon qubits: strongly gap engineered (i.e. $\delta\Delta/h\gg f_{q}$) and weakly gap engineered (i.e. $\delta\Delta/h\lesssim f_{q}$). During impact events, the weakly gap engineered qubits display correlated energy relaxation errors. Conversely, we observe no correlated $T_{1}$ decay errors on the strongly gap engineered qubits during impact events. Further, we show that the strongly gap engineered qubits are robust to significant QP poisoning induced by optical illumination of the chip, enabling us to measure qubit coherence at controlled and elevated steady-state QP densities. The robustness of this protection, coupled with the ease of achieving strong gap engineering, removes high-energy impact events as a roadblock from the path toward fault-tolerant quantum computing in superconducting qubits.
I Experimental Details
I.1 Gap engineering at the Junction
In aluminum, the superconducting gap $\Delta$ is sensitive to film thickness $t$. While $\Delta$ in thick aluminum films ($t\gtrsim 100$ nm) is close to the bulk value ($\Delta\approx\Delta_{\mathrm{bulk}}$), for thinner films ($t\lesssim 30$ nm) $\Delta$ grows quickly. We use a simple empirical model [32] for the superconducting gap $\Delta=\Delta_{\mathrm{bulk}}+A/t$, where $\Delta_{\mathrm{bulk}}=180\,\mathrm{\mu eV}$ and $A=900\,\mathrm{\mu eV{\cdot}nm}$. In Fig. 1(c), we use this model to predict the gap difference $\delta\Delta$ at the junction as a function of the thin lead thickness, holding the thick lead at 100 nm. Designing qubits with different aluminum layer thicknesses on each side of the Josephson junction permits access to different regimes of gap engineering.
The thin junction lead provides a high-gap region separating the junction from the low-gap material in the qubit capacitor. In a strongly gap engineered qubit, this barrier exceeds the energy scale of the qubit. QPs near the superconducting gap in either capacitor pad then have insufficient energy to overcome the barrier and therefore cannot tunnel across the junction. As such, strongly gap engineered qubits will suppress the QP tunneling process that produces severe qubit decay, removing the dominant impact of elevated QP density on qubit $T_{1}$. QPs should be mainly generated in the capacitor pads, and QPs in the thin layer should preferentially fall into the thick capacitor material, rapidly depleting any stray QPs in the thin layer. This should prevent any significant density reaching the junction lead, so we do not expect to need additional depositions to achieve strong suppression of QP tunneling.
I.2 Device Details
We fabricate an array of 12 transmon qubits in an all-aluminum device. We form a checkerboard placement of weakly and strongly gap engineered qubits to ensure that each class of qubit is uniformly distributed across the device, as illustrated in Fig. 1(d). The qubits are frequency-tunable with maximum frequencies in the range $f_{\text{max}}\approx 7-7.3$ GHz. We optimize the qubit geometry for coherence in this test device following Refs.  [36, 14]. Each qubit footprint is approximately 0.6 mm $\times$ 0.8 mm, with junction leads that are 2 $\mu$m long and 150 nm wide. We provide further information about the device fabrication and circuit parameters in the Supplementary Materials.
We achieve different levels of gap engineering by choosing the thicknesses of the aluminum layers at the junction. In all qubits, we choose the thick layer to be 100 nm to approach the bulk gap value $\Delta_{\mathrm{bulk}}$ and allow us to select over a wide range of gap differences $\delta\Delta$. For “weakly gap engineered” qubits, we choose 30 nm for the thin layer. This thickness targets a gap difference of around 5 GHz, substantially lower than the qubit energy for standard operating parameters. This gap difference also allows us to flux-tune qubits to explore the regimes with qubit frequencies both above and below resonance with the gap difference. For “strongly gap engineered” qubits, we choose 15 nm for the thin layer, which targets a gap difference of around 12 GHz, well above the maximum qubit frequency.
II Rapid Repetitive Correlated Sampling
Figure 2: Correlated sampling time series. Left: time series of simultaneous qubit errors summed over all qubits (top, blue), and different levels of gap engineering (GE): weakly gap engineered qubits (middle, orange) and strongly gap engineered qubits (bottom, green). Both raw data (light) and matched filtered data are shown, illustrating the location of impact events (dark). Right: corresponding histograms over simultaneous qubit errors (light) and expected rates based on independent measurements of $T_{1}$ and readout fidelities (dark). The strongly gap engineered qubits behave according to the independent error model even during impact events.
We perform rapid repetitive correlated sampling as in Ref. [3]. This involves initializing all qubits in the first excited state |$\ket{1}$, idling for 1 $\mu$s and then measuring all qubits. If the qubit is measured to be in the ground state |$\ket{0}$, we consider that a qubit error. This cycle is repeated every 100 $\mu$s over the span of 60 seconds to form a single dataset of $600,000$ samples. Over 100 datasets taken sequentially (6000 seconds total), we find 154 high-energy impact events, corresponding to an average time between events of close to 40 seconds. We include further analysis in the Supplementary Material. The corresponding event flux ($2.6\mathrm{s}^{-1}\mathrm{cm}^{-2}$) matches well to the event flux found on larger devices ($2.1\mathrm{s}^{-1}\mathrm{cm}^{-2}$ [3]). As in Ref. [3], when we perform the experiment initializing |$\ket{0}$ considering we do not find any evidence of excess rate of qubits being excited to |$\ket{1}$ during impact events.
Fig. 2 shows a single dataset featuring six events. As in Ref. [3], we use a matched filter to locate events with high confidence. Events are clearly visible on the weakly gap engineered qubits. Correspondingly, the performance of these qubits deviates noticeably from the prediction from independent $T_{1}$ and readout errors. For instance, all six weakly gap engineered qubits experience an error simultaneously two orders of magnitude more frequently than expected, indicative of highly correlated error bursts. The same events do not affect the performance of the strongly gap engineered qubits, as they continue to match the independent error model. In all 100 datasets, we find no time period where the behaviour of the strongly gap engineered qubits deviate from the independent error model, including where we see clear evidence of impacts in the weakly gap engineered qubits.
Fig. 3 shows the time series data around a single impact event. This is the event labeled $4$ shown in Fig. 2. In Fig. 3(a), we see that the event affects only the weakly gap engineered qubits, elevating their error rate significantly. By contrast, the strongly gap engineered qubits do not respond to the event, displaying error rates expected from independent $T_{1}$ and readout fidelities. During the event, we see substantial variation in error, revealing the checkerboard pattern of strongly and weakly gap engineered qubits directly in Fig. 3(b).
Figure 3: Errors during a single impact event. a) Simultaneous qubit errors in 25 ms around an impact event, summed over all qubits (blue), and different levels of gap engineering (GE): weakly gap engineered qubits (blue) and strongly gap engineered qubits (green). b) Spatial distribution of errors during the 5 ms (50 measurement cycles) immediately following the impact event (grey highlight). We include the total number of errors and the expected number of error ±$\pm 1\sigma$ from independent $T_{1}$ and readout errors. While the weakly gap engineered qubits experience greatly elevated error rates during the impact event, the strongly gap engineered qubits do not.
III Illumination of the sample
In order to further probe the qubit response to elevated QP density, we examine qubit coherence under the influence of optical illumination, which is used to introduce photons that break Cooper pairs and increase the QP density to a desired steady-state value. Specifically, we consider the density of quasiparticles as a fraction of the density of cooper pairs $x_{\mathrm{qp}}$. We use a light-emitting diode with a center wavelength of $\lambda=1200$ nm whose output is fiber coupled into the sample box, as shown in Fig. 4(a). The silicon substrate is transmissive at these optical wavelengths, and allows for the light to penetrate through the qubit and carrier chips. Unlike other works using optical illumination to implement QP injection [37], our configuration employs continuous wave flood illumination of the entire device, allowing for approximately simultaneous and uniform application of optical power flux over all qubits. We are able to vary the optical illumination power by six orders of magnitude, as well as completely disabling it, where the full-scale normalized optical power of 1.0 corresponds to 16 $\mu\mathrm{W}$ measured outside of the dilution refrigerator. Further details of the optical illumination setup can be found in the Supplementary Materials.
Fig. 4(b) shows $T_{1}$ on all the qubits at their idle frequencies, which are in the range of 6.3 GHz to 6.6 GHz, in response to varying magnitudes of optical power. As we increase the normalized optical power from 00 to $10^{-3}$, the weakly gap engineered qubits (orange circles) suffer from large $T_{1}$ reductions of over an order of magnitude. On the other hand, for the strongly gap engineered qubits (green circles), we find insignificant increase of the relaxation rates from QPs until the normalized optical power reaches $10^{-1}$. We find similar robustness in CPMG dephasing lifetimes $T_{2,\mathrm{CPMG}}$ for the strongly gap engineered qubits. The rapid reduction in qubit $T_{1}$ in the weakly gap engineered qubits limits our ability to measure $T_{2,\mathrm{CPMG}}$ during optical illumination for those qubits.
In the strongly gap engineered qubits, the preservation of coherence at high optical illumination powers allows us to carry out measurements of qubit properties in the presence of high $x_{\mathrm{qp}}$. One such measurement is the qubit frequency shift from elevated $x_{\mathrm{qp}}$. In order to account for slow qubit frequency drifts during the measurement, we use a differential Ramsey interferometry technique, where we measure the qubit frequency with no optical illumination interleaved with qubit frequency measurements at varying optical illumination powers. The interleaving period is about 10 seconds and accounts for most quasistatic drifts. Fig. 4(c) shows the qubit frequency shift $\Delta f_{q}$ of the strongly gap engineered qubits for normalized optical powers in the range of $10^{-3}$ to 1. For this range of optical powers, we observe a square-root dependence of the qubit frequency shift on the optical power. This is also the dependence between $x_{\rm qp}$ and the optical power since the magnitude of the fractional qubit frequency shift $\left|\Delta f_{q}/f_{q}\right|$ is proportional to $x_{\mathrm{qp}}$ [11, 12].
Figure 4: Measurement of qubits under illumination. a) A sketch of the illumination setup. Light from a LED is coupled into the sample box using a fiber, producing flood illumination of the device. b) Qubit coherence under illumination. Weakly gap engineered qubits (orange circles; each qubit corresponds to a shade of orange) display a reduction in $T_{1}$ beginning at very low optical powers ($1\times 10^{-4}$). Strongly gap engineered qubits display no significant change in $T_{1}$ (filled green circles; each qubit corresponds to a shade of green) or $T_{2,\mathrm{CPMG}}$ (open green circles) until high optical powers $1\times 10^{-1}$. The weakly gap engineered qubits’ response to optical power agrees with a steady-state QP density model (orange line). c) Frequency shifts of strongly gap engineered qubits under illumination. The solid line is a fit to a square-root dependence on optical power. Such a dependence is expected under a simple model where frequency shifts are proportional to QP densities and where steady-state QP density is limited by QP recombination. (Inset) Same data, plotted in linear scale.
For the weakly gap engineered qubits, we obtain the dependence of $x_{\rm qp}$ on the optical power from the qubit relaxation rates $T_{1}^{-1}(f_{q})$, measured as a function of the qubit frequency and for normalized optical powers in the range of $3.2\times 10^{-5}$ to 1. As shown in Supplementary Materials, $x_{\rm qp}$ exhibits a linear dependence on the optical power until a normalized optical power of roughly $3\times 10^{-4}$, and then $x_{\rm qp}$ exhibits a square-root dependence at larger optical powers. Notably, the inferred $x_{\mathrm{qp}}$ from the fractional qubit frequency shifts of the strongly gap engineered qubits agrees well with the inferred $x_{\mathrm{qp}}$ from the $T_{1}^{-1}(f_{q})$ spectrum of the weakly gap engineered qubits. We estimate $x_{\mathrm{qp}}\simeq 2\times 10^{-4}$ at the maximal optical illumination power, and $x_{\mathrm{qp}}\simeq 6\times 10^{-5}$ is the point where the strongly gap engineered qubits begin to show degradation (corresponding to $10^{-1}$ normalized optical power in Fig. 4(b)). Given this density exceeds what is produced even in large impact events ($x_{\mathrm{qp}}\simeq 10^{-5}$ [3]), this corroborates the insensitivity of strongly gap engineered qubits to the excess QP density induced by high-energy impact events.
IV Summary and Conclusion
We have demonstrated a gap engineering technique that removes error bursts due to high-energy impact events. Simply using a thinner aluminum junction lead, we protect our qubits from decoherence due to QP tunneling. Unlike phonon or quasiparticle trapping, this requires no additional fabrication complexity, only requiring modified deposition of aluminum.
By suppressing the mechanism of qubit decay from QP tunneling rather than directly suppressing the QP density, this technique also opens new avenues for measurement of QP-induced effects. For example, the qubit frequency shift at large QP densities was not previously resolvable given the strong suppression of $T_{1}$ in QP poisoned qubits.
That said, suppression of the QP population itself remains an important area of study. The techniques of phonon and QP trapping will be important in reducing the remaining sources of loss, including QP scattering in the capacitor material. We also note that gap engineering does not address the scrambling of two-level systems (TLSs) by high-energy impact events [4], which must be addressed either through improvements in the TLS density or via dynamic frequency-tuning to re-calibrate around the TLS scrambling in-situ. As such, rather than replacing current mitigation strategies completely, our gap engineering technique represents a substantial advance in our first line of defense, being both effective and easy to implement.
Gap engineering at the junction opens the way to strong protection from the effects of QP poisoning in large-scale devices. Future QEC experiments protected from the error floor due to impact events are now free to demonstrate the key building blocks of fault-tolerant quantum computing at scale.
Contributions
Matt McEwen, Kevin Miao, Alex Opremcak designed the experiment. Juan Atalaya, Kevin Miao provided modeling of the qubit response to QPs. Alex Opremcak, Alex Crook, Jenna Bovaird, John Mark Kreikebaum contributed to the design and fabrication of the experimental device. Kevin Miao, Alex Opremcak, Alex Bilmes, Nick Zobrist, Evan Jeffrey constructed the experimental apparatus. Kevin Miao, Matt McEwen, Alex Bilmes calibrated and measured the experimental device. Bicheng Ying, Matt McEwen contributed key software infrastructure for performing measurements. Alex Bilmes, Andreas Bengtsson, Alex Crook, Hung-Shen Chang, Andrew Dunsworth conducted measurements of test devices. Julian Kelly, Yaxing Zhang provided modeling of limiting loss mechanisms. Ebrahim Forati provided electromagnetic simulations of devices. Rajeev Acharya, Justin Iveland, Wayne Liu, Seon Kim, and Brian Burkett contributed to the fabrication, characterization, and processing techniques used in the devices. Matt McEwen, Kevin Miao, Alex Opremcak compiled the manuscript. All authors contributed to revising the manuscript.
Acknowledgements
We thank Lev Ioffe and Lara Faoro for helpful discussions. We thank Alexander Korotkov for helpful comments on the manuscript. We thank the entire Google Quantum AI team for maintaining the hardware, software, cryogenics, and electronics infrastructure that enabled this experiment, and for producing an environment where this work is possible.